OpenStudy (anonymous):

What polynomial has roots of –1, –2, and 3 ?

OpenStudy (callisto):

(x+1)(x+2)(x-3) = (x^2 + 3x+2) (x-3) = x^3 -x-6

OpenStudy (anonymous):

(x+1)(x+2)(x-3) = (x^3+3x^2+2x)+(-3x^2-9x-6) = x^3-7x-6

OpenStudy (anonymous):

shouldnt it be -7x?

OpenStudy (anonymous):

Callisto, I believe you screwed up on your 'x' terms, let me double check my math too

OpenStudy (anonymous):

oh nvm. ok

OpenStudy (anonymous):

Yes, my answer appears to be correct

OpenStudy (callisto):

sorry my fault :(